OpenStudy (adrianabell):

how many solution are there to the equation 5(x+10)-25=5x+25

OpenStudy (linktra):

Because this equation is has equal sides, and also has only 1 variable, you can test how many solutions it has by doing one where x is a negative number, and one where x is a positive number. \[5(x+10)-25=5x+25\] First I will do where x is positive. I will use x=3. \[5(3+10)-25=5(3)+25\] Now I will complete what is in parenthesis. \[5(13)-25=5(3)+25\] Now I will do the multiplication of both sides. \[65-25=15+25\] Now I will add and subtract on both sides. \[40=40\] So now that I know this equation works with the positive value of x, I will work out the equation with a negative value for x. First write the original equation. \[5(x+10)-25=5x+25\] Now replace x with the negative value. I will use x=-4. \[5(-4+10)-25=5(-4)+25\] Now work out the parentheses. \[5(6)-25=5(-4)+25\] Now multiply on both sides. \[30-25=-20+25\] Finally add and subtract on both sides. \[5=5\] Because this equation is true with both a positive value and a negative value for x, it means that this equation has an infinite number of solutions.

OpenStudy (eliesaab):

It reduces 5 (x + 10) - 25 = 5 x + 25 5 x +50 -25 =5x +25 0 x =0 Which is true for any x